OpenStudy (anonymous):

Find the points on the given curve where the tangent line is horizontal or vertical r=f(q)=cos(q)+sin(q)

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=polar+plot+r+%3D+sin+theta+%2B+cos+theta it's circle ... you can do this using rectangular coorinates

OpenStudy (experimentx):

If you want to do it in polar coordinates then, http://tutorial.math.lamar.edu/Classes/CalcII/PolarTangents.aspx

OpenStudy (anonymous):

got it Thanks dude

OpenStudy (experimentx):

it seems that equation of line in polar coordinate is \[ y = mx + c \implies r(\sin\theta - m\cos \theta) = c \implies r = \frac{c}{(\sin\theta - m\cos\theta)}\] Find the values of m, using the above formula,

OpenStudy (experimentx):

I mean from the link i posted above, and equate it with 0 and \( \infty \) and find the value of \(r\) and \( \theta \) and put them in above equation of line to find c, then you have your tangents.